OpenStudy (anonymous):

Solve using completing the square and quadratic formula. 3y^2+4y+2=0

OpenStudy (callisto):

using quad. formula, y=[-4±√(4^2-4*3*2)]/2(3) = (-2±√2i)/3

OpenStudy (anonymous):

y^2+4y/3 + 2/3 =0 so (y+2/3)^2 -4/9 +2/3 =0 So (y+2/3) = sqrt(-2/9) etc....